# Application-Driven Innovation in Machine Learning

### v1

@@ Abstract
As applications of machine learning proliferate, innovative algorithms inspired by specific real-world challenges have become increasingly important. Such work offers the potential for significant impact not merely in domains of application but also in machine learning itself. In this paper, we describe the paradigm of application-driven research in machine learning, contrasting it with the more standard paradigm of methods-driven research. We illustrate the benefits of application-driven machine learning and how this approach can productively synergize with methods-driven work. Despite these benefits, we find that reviewing, hiring, and teaching practices in machine learning often hold back application-driven innovation. We outline how these processes may be improved.
@@ 1 Introduction
Machine learning (ML) is increasingly being used across diverse fields and sectors, with significant impacts for society. ML is being used in healthcare to analyze genetic markers, process medical imagery, and digitize health records [16]. ML is being used in climate science to speed up physical simulations, parse satellite data, and forecast extreme events [31, 40]. ML is being used in heavy industry to control complex processes, optimize supply chains, and design new materials [17]. These kinds of applications, and many others [51], involve diverse ML tools being used in a multiplicity of ways.
The widespread use of machine learning across society draws on decades of innovation in core ML algorithms, but also on another type of ML research: Application-driven innovation. ML algorithms designed in blue-sky, methods-focused research continue to fall short when used directly for applications. Bridging the gap requires thoughtful consideration of the challenges of real-world tasks and the properties of real-world data, and welcoming use cases into the research process. This approach to ML research has much to contribute to broader innovation in ML methods, as well as downstream applications. However, application-driven innovation has characteristics that have led too often to it being under-valued within the ML community or perceived as out-of-scope [42].
@@ 2.1 Methods-driven research
The mainstream paradigm of innovation in machine learning, which may be referred to as methods-driven research, focuses on the identification of algorithms with certain target properties and evaluates these properties using datasets. Methods-driven ML research has followed various trends over the past several decades (see also [50]). For illustration and comparison, we review key aspects of the paradigm as it currently exists, especially in deep learning.
@@
    Stereotyped benchmarks. Many of the innovations celebrated within the ML community arise through experiments on well-established benchmarks such as ImageNet [9], MS COCO [24], the OpenAI Gym [7], etc. While new benchmarks and variations of existing ones continue to be introduced, the number of fundamentally distinct problems remains very limited and the most popular benchmarks have largely retained their hegemony. For example, the three datasets and testbeds above were cited a total of 26,000 times in 2023.
@@
    Massive datasets. Algorithms are assumed to benefit from more data. The subtleties of the data and how it was collected are generally considered to be less important than the quantity [2], despite mounting evidence that data composition is key [14]. For example, in creating ImageNet and the training datasets for most large language models, researchers scraped large amounts of raw Internet data, with annotations provided at scale by teams of generally untrained annotators [9]. Small datasets are generally perceived as less important or trustworthy in assessing an algorithm’s effectiveness.
@@
    Problem-agnostic methods. As with metrics for evaluation, the algorithms utilized are designed to be as general as possible. It is common to use a variety of very different datasets to evaluate a new method. It is also increasingly common, with the rise in foundation models [6], to design an algorithm to simultaneously perform a range of different tasks and expect it to generalize well to other tasks.
It is often taken for granted within the ML community that these principles are the only way to produce meaningful innovation in machine learning. Yet as many authors have detailed [4, 15], these are not inevitabilities but rather are implicit choices that reflect the values and priorities within the field.
@@ 2.2 Application-driven research
Machine learning researchers are frequently shocked to discover that algorithms designed using methods-driven innovation fall short in many real-world use cases. For example, in many remote sensing problems, a simple random forest approach proves more effective than highly regarded computer vision approaches such as Vision Transformers [13]. Such gaps between expectation and reality (which indeed are only apparent when ML researchers engage with end users) illustrate that methods-driven research can be usefully complemented by other types of innovation.
@@
    Real-world tasks. Where methods-driven research evaluates success based on standardized benchmarks, application-driven ML concentrates on a real-world problem, or a family of related problems. The goal is to design an algorithm that satisfies the specific needs of this problem, based on how the ML will ultimately be used. Framing the problem as a machine learning task, or set of tasks, is often part of the challenge. An increasing number of application-specific benchmarks, such as iNaturalist [47], ClimSim [53], WILDS [21], MIMIC-CXR [18], CityLearn [48], and Grid2Op [28], have been created, often with considerable time and effort. These capture different data and task properties than standard ML benchmarks and are extremely useful for ADML work. However, no single fixed set of benchmarks is sufficient to capture the full scope of real-world uses for ML, and trying to fit applications into the standard paradigm of ML is limiting.
@@
    Application-specific evaluation metrics. One of the key reasons that methods-driven ML fails to be impactful in some practical applications is that the set of appropriate evaluation metrics varies greatly between real-world use cases. For example, labeling land cover types from satellite imagery is a common challenge in remote sensing, and while accuracy matters, the computational cost is also a major factor since algorithms must scale over large areas, often on limited infrastructure [39]. By contrast, identification of animals from photographs is an important challenge in biodiversity monitoring, and here uncertainty quantification can be essential so that statistics can be appropriately incorporated into ecological models [49]. From a method-driven ML perspective, both of these tasks are image classification, but for ADML, the downstream criteria for success are different, so evaluation must include these metrics as well as metrics such as test loss and accuracy.
    Standard ML evaluation procedures and frameworks may indeed not apply at all. For example, standard uniform-at-random assignment of instances to test splits can drastically over-estimate performance of a model used in conditions or regions unseen during training [34, 30]; blocked, clustered, or buffered holdout sets are more appropriate to assess performance in out-of-sample use cases for prediction domains with spatial, temporal, or other hierarchical structures [38, 22]. It is also worth noting that common goals such as interpretability, robustness, and generalization often have multiple nuanced definitions depending on the particular application. The adversarial robustness metrics commonly used in ML, for example, represent a distinct notion from the Lyapunov stability guarantees expected in certain engineered systems (see e.g. [10]). In applications of reinforcement learning, sample efficiency, sensitivity to hyperparameter choices (because there is no analogy of cross validation), and stability of the deployed controller are critical, whereas method-driven evaluation in reinforcement learning assumes unfettered access to simulator data required to tune ever increasingly complex agents for peak performance.
@@
    Auxiliary domain knowledge. While methods-driven ML typically focuses on datasets that are as large as possible, application-driven ML often does not have this luxury. While large datasets remain very useful if they are available, often real-world problems place more data (especially more labeled data) at a premium. In cases where datasets are limited, ADML approaches instead often focus on incorporating as much domain knowledge and auxiliary information as possible, as well as careful selection/curation of labeled data. By contrast, methods-driven research often is unwilling or unable to incorporate such problem-specific information, since distinct domain-specific challenges may be put in a single bucket. For example, if computer vision models pretrained on ImageNet are used on remote sensing data, they must simplify the wide array of sensors used to collect such data, which extend beyond RGB channels. By contrast, ADML approaches for remote sensing explicitly incorporate such diverse inputs [44, 46]. Similarly, location information from datapoints is increasingly being incorporated into ADML approaches in areas such as agriculture and ecology [27, 45, 20].
@@
    Problem-informed methods. Related to the above point on auxiliary data is how the structure of the data and the problem itself is reflected in the choice of algorithm. At the most basic level, many machine learning approaches are chosen based upon a high-level task classification – thus, convolutional neural networks may be used for image-structured data, LSTMs for time series, etc. However, the structure of a particular problem may be extended beyond such coarse categories. The domain of physics-informed ML often intersects with ADML, if the underlying physics of a problem is considered – rules such as conservation of energy or mass, or known differential equations governing the variables [19]. Similarly, ML methods for engineered systems often should incorporate known constraints between variables [11]. In cases where ML methods must come with provable guarantees (e.g., in safety-critical systems), incorporating some domain knowledge may in fact be necessary, in order to enable theoretical analysis [12]. Constraints or rules can often be incorporated into the design of the algorithm, for example, via custom neural network layers or a special loss function [19]. Often, input data can be represented in a helpful way with expert knowledge from the application domain, as when climatological variables can be considered as a graph and processed using graph neural networks [43].
@@ 2.3 Contributions to machine learning
Paradoxically, while ADML research is problem-centric, this approach has not merely improved the performance of ML in specific use cases, but has done much to advance ML research as a whole. Methods originally tailored for a specific problem have often proven useful to a variety of seemingly dissimilar problems. For example, Fourier Neural Operators [23], originally developed to solve the differential equations governing fluid flow, have now been used in applications as diverse as climate data superresolution and materials property prediction [52, 37]. This is counterintuitive, but reflects the fact that real-world challenges may share more underlying structure with each other than with the often artificial benchmarks used in methods-driven research.
The use cases raised in ADML can also tie in to broad challenges considered in methods-driven research, or may raise new challenges that can be abstracted away from the task at hand. For example, the iNaturalist benchmark was originally constructed as a byproduct of computer vision research on a real-world biodiversity problem [47], but it has now become a popular test case for learning very long-tailed distributions. Many problems in ML for climate and weather modeling pose a special type of nonstationarity challenge, since climate change means that training data drawn from past observations will not necessarily reflect patterns in future data [31].
@@ 3 Reviewing
Of course, not every paper on applications of ML is well-suited to ICML. Many such papers are most appropriate to publication in the domain of application, or in a cross-disciplinary venue such as Environmental Data Science or Nature Digital Medicine. These papers may, for example, be focused on applications of existing techniques to help address a narrow problem in the domain of interest. However, other papers represent ML contributions – either in the methodology or in the kind of task being studied. A methodology contribution may be developed for one application but end up being broadly relevant across ML. A task-related contribution may provide a useful testbed for evaluating ML according to real-world challenges, or highlight metrics and considerations not previously regarded as important to ML evaluation. Both types of contributions thus have much to offer the ML community and fall squarely within the scope of ML venues.
@@
  • Unfamiliar benchmarks. Reviewers are often confused when a paper does not test against the suite of benchmarks that have become standard in methods-driven ML. Often, ADML papers do not use such benchmarks because the structure of standard datasets, or the goals of standard benchmarking tasks, differ significantly from the real-world data under consideration. For example, image classification algorithms developed for remote sensing are generally not ideal for ImageNet and vice versa. Researchers may also avoid certain standardized benchmarks because of toxic content [35, 5], ethics issues involved in their creation [1, 33], known inaccuracies or biases [26, 8]. Despite these considerations, reviewers frequently call for experiments on the “stereotypical” datasets if they are not familiar with the datasets used in the paper, even when such datasets are high-quality, published, and supported by appropriate baselines.
@@
  • Too simple. The common criticism that a novel method is “too simple” is often used when the reviewer is surprised that it outperforms complicated “SoTA” algorithms. ML scholarship often prizes the appearance of complexity, which can lead to simple insights being disregarded where ideally they would be prized [50, 25]. By contrast, ADML researchers often have an especial focus on usability and scalability, so the lack of unnecessary complexity can indeed be a goal, and may be the result of distilling a complex method down to its core innovation.
@@
These failure modes of the review process do not necessarily arise from the high-level reviewing guidelines of ML venues, but rather are the consequence of interpreting these guidelines according to the methods-driven research approach detailed above. Below, we attempt to provide guidance to reviewers and area chairs on how to interpret common rubrics for ML research excellence in the case of application-driven work. These may also be useful to ML venues interested in specifying reviewing guidelines with a greater degree of granularity.
@@
  • Clarity. A common failure of ADML papers is writing for the wrong audience. For publication in an ML venue, a paper should provide background in the relevant application domain and dataset(s), definitions of any jargon used, and clear motivation for the problem at hand. The authors should, of course, also provide methodological details to a higher standard than that sometimes expected in applied venues.
@@
Overall, we are seeing slow but meaningful progress in ADML publication opportunities. In addition to the growing pool of researchers empowered within the ML community, several ML venues have created focused tracks aimed at specific kinds of ADML publications. The NeurIPS Datasets and Benchmarks track has become a home for papers on innovative application-aligned tasks, while the JMLR special issue on climate change and AAAI Social Impact track have similarly broadened the scope of available venues. We hope, however, that ultimately some such workarounds become obsolete as ADML research becomes better recognized and supported as part of mainstream ML innovation.
Many ADML researchers have gravitated towards subcommunities of the broader field of ML, which present opportunities for exchange with like-minded individuals. These spaces include workshop series such as ML4D and “Tackling Climate Change with Machine Learning”, and conferences such as ML4H and Climate Informatics. These spaces serve an important purpose complementary to ML venues such as ICML and NeurIPS, furthering discourse between researchers in ML and specific application domains, and in some cases serving as publication venues. However, they are not a replacement for the recognition of ADML research within the wider ML community, and the relegation of such work to smaller, more narrowly focused spaces stands to silo ADML works that would be useful and interesting to ML research at large.

### v2

# Position: Application-Driven Innovation in Machine Learning

@@ Abstract
In this position paper, we argue that application-driven research has been systemically under-valued in the machine learning community. As applications of machine learning proliferate, innovative algorithms inspired by specific real-world challenges have become increasingly important. Such work offers the potential for significant impact not merely in domains of application but also in machine learning itself. In this paper, we describe the paradigm of application-driven research in machine learning, contrasting it with the more standard paradigm of methods-driven research. We illustrate the benefits of application-driven machine learning and how this approach can productively synergize with methods-driven work. Despite these benefits, we find that reviewing, hiring, and teaching practices in machine learning often hold back application-driven innovation. We outline how these processes may be improved.
applied ML, AI for Good, AI for Science, health, climate
@@ 1 Introduction
Machine learning (ML) is increasingly being used across diverse fields and sectors, with significant impacts for society. ML is being used in healthcare to analyze genetic markers, process medical imagery, and digitize health records (ghassemi2020review). ML is being used in climate science to speed up physical simulations, parse satellite data, and forecast extreme events (monteleoni2013climate; rolnick2022tackling). ML is being used in heavy industry to control complex processes, optimize supply chains, and design new materials (gopaluni2020modern). These kinds of applications, and many others (wang2023scientific), involve diverse ML tools being used in a multiplicity of ways.
The widespread use of machine learning across society draws on decades of innovation in core ML algorithms, but also on another type of ML research: Application-driven innovation. ML algorithms designed in blue-sky, methods-focused research continue to fall short when used directly for applications. Bridging the gap requires thoughtful consideration of the challenges of real-world tasks and the properties of real-world data, and the welcoming of use cases into the research process. This approach to ML research has much to contribute to broader innovation in ML methods, as well as downstream applications. However, application-driven innovation has characteristics that have led too often to it being under-valued within the ML community or perceived as out-of-scope (rudin2014machine).
@@ 2.1 Methods-Driven Research
The mainstream paradigm of innovation in machine learning methods, which may be referred to as methods-driven research, focuses on the identification of algorithms with certain target properties and evaluates these properties using datasets. Methods-driven ML research has followed various trends over the past several decades (see also wagstaff2012machine). For illustration and comparison, we review key aspects of the paradigm as it currently exists, especially in deep learning.
@@
    Stereotyped benchmarks. Many of the innovations celebrated within the ML community arise through experiments on well-established benchmarks such as ImageNet (deng2009imagenet), MS COCO (lin2014microsoft), the OpenAI Gym (brockman2016openai), etc. While new benchmarks and variations of existing ones continue to be introduced, the number of fundamentally distinct problems remains very limited and the most popular benchmarks have largely retained their hegemony. For example, the three datasets and testbeds above were cited a total of 26,000 times in 2023.
@@
    Massive datasets. Algorithms are assumed to benefit from more data. The subtleties of the data and how it was collected are generally considered to be less important than the quantity (bender2021dangers), despite mounting evidence that data composition is key (fang2022data). For example, in creating ImageNet and the training datasets for most large language models, researchers scraped large amounts of raw Internet data, with annotations provided at scale by teams of generally untrained annotators (deng2009imagenet). Small datasets are generally perceived as less important or trustworthy in assessing an algorithm’s effectiveness.
@@
    Problem-agnostic methods. As with metrics for evaluation, the algorithms utilized are designed to be as general as possible. It is common to use a variety of very different datasets to evaluate a new method. It is also increasingly common, with the rise in foundation models (bommasani2021opportunities), to design an algorithm to simultaneously perform a range of different tasks and expect it to generalize well to other tasks.
It is often taken for granted within the ML community that these principles are the only way to produce meaningful innovation in machine learning. Yet as many authors have detailed (birhane2022values; ghassemi2022machine), these are not inevitabilities but rather are implicit choices that reflect the values and priorities within the field.
@@ 2.2 Application-Driven Research
Machine learning researchers are frequently shocked to discover that algorithms designed using methods-driven innovation fall short in many real-world use cases. For example, in many remote sensing problems, a simple random forest approach proves more effective than highly regarded computer vision approaches such as Vision Transformers (dosovitskiy2020image). Such gaps between expectation and reality (which indeed are only apparent when ML researchers engage with end users) illustrate that methods-driven research can be usefully complemented by other types of innovation.
@@
    Real-world tasks. Where methods-driven research evaluates success based on standardized benchmarks, application-driven ML concentrates on a real-world problem, or a family of related problems. The goal is to design an algorithm that satisfies the specific needs of this problem, based on how the ML will ultimately be used. Framing the problem as a machine learning task, or set of tasks, is often part of the challenge. An increasing number of application-specific benchmarks, such as iNaturalist (van2018inaturalist), ClimSim (yu2023climsim), WILDS (koh2020wilds), MIMIC-CXR (johnson2019mimic), CityLearn (vazquez2020citylearn), and Grid2Op (marot2021learning), have been created, often with considerable time and effort. These capture different data and task properties than many standard ML benchmarks and are extremely useful for ADML work. However, no single fixed set of benchmarks is sufficient to capture the full scope of real-world uses for ML.
@@
    Application-specific evaluation metrics. One of the key reasons that methods-driven ML fails to be impactful in some practical applications is that the set of appropriate evaluation metrics varies greatly between real-world use cases. For example, labeling land cover types from satellite imagery is a common challenge in remote sensing, and while accuracy matters, the computational cost is also a major factor since algorithms must scale over large areas, often on limited computational infrastructure (robinson2019large). Identification of animals from photographs is an important challenge in biodiversity monitoring, and here uncertainty quantification can be essential so that statistics can be appropriately incorporated into ecological models (villon2020new). From a method-driven ML perspective, both of these tasks are image classification, but for ADML, the downstream criteria for success are different, so evaluation must include these metrics as well as metrics such as test loss and accuracy.
    In some cases, standard ML evaluation procedures and frameworks may indeed not apply at all. For example, standard uniform-at-random assignment of instances to test splits can drastically over-estimate performance of a model used in conditions or regions unseen during training (ploton2020spatial; meyer2022machine); blocked, clustered, or buffered holdout sets are more appropriate to assess performance in out-of-sample use cases for prediction domains with spatial, temporal, or other hierarchical structures (roberts2017cross; le2014spatial). It is also worth noting that common goals such as interpretability, robustness, and generalization often have multiple nuanced definitions depending on the particular application. The adversarial robustness metrics commonly used in ML, for example, represent a distinct notion from the Lyapunov stability guarantees expected in certain engineered systems (see e.g. donti2022climate). In applications of reinforcement learning, sample efficiency, sensitivity to hyperparameter choices (because there is no analogy of cross validation), and stability of the deployed controller are especially critical, whereas method-driven evaluation in reinforcement learning often assumes unfettered access to simulator data required to tune ever increasingly complex agents for peak performance.
@@
    Auxiliary domain knowledge. While methods-driven ML typically focuses on datasets that are as large as possible, application-driven ML often does not have this luxury. While large datasets remain very useful if they are available, often real-world problems place more data (especially more labeled data) at a premium. In cases where datasets are limited, ADML approaches instead often focus on incorporating as much domain knowledge and auxiliary information as possible, as well as careful selection/curation of labeled data. By contrast, methods-driven research often is unwilling or unable to incorporate such problem-specific information, since distinct domain-specific challenges may be put in a single bucket. For example, if computer vision models pretrained on ImageNet are used on remote sensing data, they must simplify the wide array of sensors used to collect such data, which extend beyond RGB channels. By contrast, ADML approaches for remote sensing explicitly incorporate such diverse inputs (russwurm2024meta; tseng2023lightweight). Similarly, location information from datapoints is increasingly being incorporated into ADML approaches in areas such as agriculture and ecology (mac2019presence; tseng2022timl; klemmer2023satclip).
@@
    Problem-informed methods. Related to the above point on auxiliary data is how the structure of the data and the problem itself is reflected in the choice of algorithm. At the most basic level, many machine learning approaches are chosen based upon a high-level task classification – thus, convolutional neural networks may be used for image-structured data, LSTMs for time series, etc. However, the structure of a particular problem may be extended beyond such coarse categories. The domain of physics-informed ML often intersects with ADML, if the underlying physics of a problem is considered – e.g., rules such as conservation of energy or mass, or known differential equations governing the variables (kashinath2021physics). Similarly, ML methods for engineered systems often should incorporate known constraints between variables (donti2021machine). In cases where ML methods must come with provable guarantees (e.g., in safety-critical systems), incorporating some domain knowledge may in fact be necessary, in order to enable theoretical analysis (donti2020enforcing). Constraints or rules can often be incorporated into the design of the algorithm, for example, via custom neural network layers or a special loss function (kashinath2021physics). Desirable objectives can also sometimes be encoded into a loss function, as in the case of optimizing limited resources in public health (wang2023scalable; mate2022field). Often, input data can be represented in a helpful way with expert knowledge from the application domain, as when climatological variables can be considered as a graph and processed using graph neural networks (cachay2020graph).
@@
Note that our discussions of both methods-driven and application-driven research refer to works introducing machine learning methodology, rather than those focused on theory. Of course, many papers include theoretical analysis alongside a largely empirical evaluation, and it is possible for such works to be either methods-driven or application-driven, since theory can be valuable in either case. As an example of theoretical analysis in an ADML work, see duval2023faenet, which introduces novel algorithms for materials science applications. Here, the theoretical results justify the property of equivariance under group transformations, which is important for the applications in question.
@@ 2.3 Contributions to Machine Learning
Paradoxically, while ADML research is problem-centric, this approach has not merely improved the performance of ML in specific use cases, but has done much to advance ML research as a whole. Methods originally tailored for a specific problem have often proven useful to a variety of seemingly dissimilar problems. For example, Fourier Neural Operators (li2020fourier), originally developed to solve the differential equations governing fluid flow, have now been used in applications as diverse as climate data superresolution and materials property prediction (yang2023fourier; rashid2022learning). This is counterintuitive, but reflects the fact that real-world challenges may share more underlying structure with each other than with the often artificial benchmarks used in methods-driven research.
The use cases raised in ADML can also tie in to broad challenges considered in methods-driven research, or may raise new challenges that can be abstracted away from the task at hand. For example, the iNaturalist benchmark was originally constructed as a byproduct of computer vision research on a real-world biodiversity problem (van2018inaturalist), but it has now become a popular test case for learning very long-tailed distributions. Many problems in ML for climate and weather modeling pose a special type of nonstationarity challenge, since climate change means that training data drawn from past observations will not necessarily reflect patterns in future data (monteleoni2013climate).
@@ 3 Reviewing
Of course, not every paper on applications of ML is well-suited to ICML. Many such papers are most appropriate to publication in the domain of application, or in a cross-disciplinary venue such as Environmental Data Science or Nature Digital Medicine. These papers may, for example, be focused on applications of existing techniques to help address a narrow problem in the domain of interest. However, other papers represent ML contributions – either in the methodology or in the kind of task being studied. A methodology contribution may be developed for one application but end up being broadly relevant across ML. A task-related contribution may provide a useful testbed for evaluating ML according to real-world challenges, or highlight metrics and considerations not previously regarded as important to ML evaluation. Both types of contributions thus have much to offer the ML community and fall squarely within the scope of venues such as ICML. Examples of past works of this form which were nevertheless published outside of mainstream ML venues include ronneberger2015u; raissi2019physics; kurth2023fourcastnet; chen2021enforcing.
@@
  • Unfamiliar benchmarks. Reviewers are often confused when a paper does not test against the suite of benchmarks that have become standard in methods-driven ML. Often, ADML papers do not use such benchmarks because the structure of standard datasets, or the goals of standard benchmarking tasks, differ significantly from the real-world data under consideration. For example, image classification algorithms developed for remote sensing are generally not ideal for ImageNet and vice versa. Researchers may also avoid certain standardized benchmarks because of toxic content (prabhu2020large; birhane2021multimodal), ethics issues involved in their creation (asano2021pass; perrigo2023exclusive), or known inaccuracies or biases (luccioni2023bugs; crawford2021excavating). Despite these considerations, reviewers frequently call for experiments on the “stereotypical” datasets if they are not familiar with the datasets used in the paper, even when such datasets are high-quality, published, and supported by appropriate baselines.
@@
  • Too simple. The common criticism that a novel method is “too simple” is often used when the reviewer is surprised that it outperforms complicated “SoTA” algorithms. ML scholarship often prizes the appearance of complexity, which can lead to simple insights being disregarded where ideally they would be prized (wagstaff2012machine; lipton2018troubling). By contrast, ADML researchers often have an especial focus on usability and scalability, so the lack of unnecessary complexity can indeed be a goal, and may be the result of distilling a complex method down to its core innovation.
@@
These failure modes of the review process do not necessarily arise from the high-level reviewing guidelines of ML venues, but rather are the consequence of interpreting these guidelines according to the methods-driven research approach detailed above. Below, we attempt to provide guidance to reviewers and area chairs on how to interpret common rubrics for ML research excellence in the case of application-driven work. These may additionally be useful to ML venues interested in specifying reviewing guidelines with a greater degree of granularity. (See also the reviewer guidelines for the AAAI Special Track on AI for Social Impact.)
@@
  • Clarity. A common failure of ADML papers is writing for the wrong audience. For publication in an ML venue, a paper should provide background in the relevant application domain and dataset(s), definitions of any jargon used, and clear motivation for the problem at hand. The authors should, of course, also provide ML methodological details to a higher standard than that sometimes expected in applied venues.
@@
Overall, we are seeing slow but meaningful progress in ADML publication opportunities. In addition to the growing pool of researchers empowered within the ML community, several ML venues have created focused tracks aimed at specific kinds of ADML publications. The NeurIPS Datasets and Benchmarks track has become a home for papers on innovative application-aligned tasks, while the JMLR special issue on climate change, AAAI Social Impact track, and IAAI conference have similarly broadened the scope of available venues. The further development of special tracks oriented towards application-driven work could be a valuable next step for ML publication venues. We hope, however, that ultimately such workarounds become obsolete as ADML research becomes better recognized and supported as part of mainstream ML innovation.
Many ADML researchers have gravitated towards subcommunities of the broader field of ML, which present opportunities for exchange with like-minded individuals. These spaces include workshop series such as ML4D, “Tackling Climate Change with Machine Learning”, and “RL for Real Life”, and conferences such as ML4H and Climate Informatics. These spaces serve an important purpose complementary to ML venues such as ICML and NeurIPS, furthering discourse between researchers in ML and specific application domains, and in some cases serving as publication venues. However, they are not a replacement for the recognition of ADML research within the wider ML community, and the relegation of such work to smaller, more narrowly focused spaces stands to silo ADML works that would be useful and interesting to ML research at large.